# Splittings in a complete local ring and decomposition its group of units

### v1

@@ 3. The general case
Let $S=\{x_{i}:i\in I\}\subseteq k$ be a transcendence basis of the extension $k/\mathbb{Q}$. Since the set $S$ is algebraically independent over $\mathbb{Q}$, the canonical morphism of $\mathbb{Q}$-algebras from the polynomial ring $\mathbb{Q}[X_{i}:i\in I]$ to $k$ given by $X_{i}\mapsto x_{i}$ is injective. Let $R=\mathbb{Q}[x_{i}:i\in I]$ be the image of this ring map. Since $f$ is surjective, for each $i$, there exists some $a_{i}\in A$ such that $f(a_{i})=x_{i}$. Then again by the universal property of polynomial rings, we have a (unique) morphism of $\mathbb{Q}$-algebras $R\rightarrow A$ which maps each $x_{i}$ into $a_{i}$. To extend this map to $k_{0}=\mathbb{Q}(x_{i}:i\in I)$, the field of fractions of the integral domain $R$, we need to show that the image of every nonzero element $P(x_{i_{1}},\ldots,x_{i_{n}})$ of $R$ under this map is invertible in $A$. Note that $A$ is a local ring with the maximal ideal $M=\operatorname{Ker}(f)$, because $M$ is nilpotent. So it suffices to show that $P(a_{i_{1}},\ldots,a_{i_{n}})\notin M$. Since $f$ is a $\mathbb{Q}$-algebra map, we have $f(P(a_{i_{1}},\ldots,a_{i_{n}}))=P(x_{i_{1}},\ldots,x_{i_{n}})\neq 0$. This shows that $P(a_{i_{1}},\ldots,a_{i_{n}})$ is invertible in $A$. So we can extend the above ring map to obtain a ring map $h:k_{0}\rightarrow A$ such that $fh:k_{0}\rightarrow k$ is the inclusion map. This, in particular, shows that $A$ is a $k_{0}$-algebra and $h$ is a $k_{0}$-algebra morphism. To extend this ring map $h$ to $k$, since $k$ is an algebraic extension over $k_{0}$, then by Corollary 3.9, the extension $k/k_{0}$ is formally étale. This observation allows us to show that the natural map $\operatorname{Hom}_{k_{0}}(k,A)\rightarrow\operatorname{Hom}_{k_{0}}(k,k)$ given by $g^{\prime}\mapsto fg^{\prime}$ is bijective where the $\operatorname{Hom}_{k_{0}}(-,-)$ is computed in the category of $k_{0}$-algebras. Indeed, if $h^{\prime}:k\rightarrow k$ is a morphism of $k_{0}$-algebras then we have the following commutative diagram of rings:
@@
Remark 3.11.
Here we will provide an alternative proof for the last part of the proof of the above lemma (without using Corollary 3.9): we can write $k$ as the filtered union (direct limit) of all finite subextensions of $k_{0}$, i.e., $k=\bigcup_{\alpha}k_{\alpha}=\lim\limits_{\overrightarrow{\alpha}}k_{\alpha}$ where each subfield $k_{\alpha}$ of $k$ is a finite extension of $k_{0}$. But in the characteristic zero, every (finite) algebraic extension of fields is separable. We also know that every finite separable extension of fields is formally étale. Thus for any $\alpha$, the field extension $k_{0}\subseteq k_{\alpha}$ is formally étale, and so the natural map $\phi_{\alpha}:\operatorname{Hom}_{k_{0}}(k_{\alpha},A)\rightarrow\operatorname{Hom}_{k_{0}}(k_{\alpha},k)$ is bijective. Then taking inverse limits, the induced map $\lim\limits_{\overleftarrow{\alpha}}\phi_{\alpha}:\lim\limits_{\overleftarrow{\alpha}}\operatorname{Hom}_{k_{0}}(k_{\alpha},A)\rightarrow\lim\limits_{\overleftarrow{\alpha}}\operatorname{Hom}_{k_{0}}(k_{\alpha},k)$ is also bijective. This shows that the natural map:
$\phi:\operatorname{Hom}_{k_{0}}(k,A)=\operatorname{Hom}_{k_{0}}(\lim\limits_{\overrightarrow{\alpha}}k_{\alpha},A)\rightarrow\operatorname{Hom}_{k_{0}}(k,k)$
given by $g^{\prime}\mapsto fg^{\prime}$ is bijective, because this is due to the commutativity of the following diagram of sets:
where the vertical arrows are the canonical isomorphisms (noting that the Hom functor converts direct limits in the first factor to inverse limits, see Remark 3.2).
@@
Consider the short exact sequence:
@@
where the second map is multiplication by $r$ on the $\operatorname{Ext}_{R}^{d}(M,N)$. But by Stacks Project [2, Tag 00LV], this map vanishes since $rN=0$, then in either case of the hypothesis, making $\operatorname{Ext}_{R}^{d}(M,N)$ vanish by exactness. ∎
@@
The above corollary is quite sufficient for our purposes. However, we should add that this result holds more generally: Let $R$ be a regular Noetherian ring of Krull dimension $n$ and $M,N$ be modules over $R$. If there is an element $r\in R$ such that $r$ is regular on $M$ and $rN=0$, then $\operatorname{Ext}_{R}^{n}(M,N)=0$. In fact, this is proved by the same argument as the above by applying a theorem of Serre [9] which asserts that if $R$ is regular Noetherian of Krull dimension $n$, then for any two $R$-modules $M,N$ one has $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>n$. The above corollary also holds over a Dedekind domain as well.
Recall that by a complete local ring we mean a local ring $(R,M)$ such that the canonical ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ is an isomorphism. Note that in this definition $R$ is not assumed to be Noetherian (contrary to the classical literature, see e.g. [1], where complete local rings were also assumed to be Noetherian). It is interesting to note that the completion of every ring at a maximal ideal is always a complete local ring. More precisely, if $M$ is a maximal ideal of a ring $R$, then it can be easily seen that $\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ is a complete local ring with the residue field $R/M$.
@@
$\mathbf{(ii)}$ If $\operatorname{char}(k)=p>0$, then the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
@@
given by $g\mapsto\pi g$ is surjective for all $n\geqslant 1$, where $\pi:R/M^{n+1}\rightarrow R/M^{n}$ is the canonical ring map and the $\operatorname{Hom}(-,-)$ is computed in the category of commutative rings. Suppose we have proven this, then using Remarks 3.2 and 3.15, the natural map:
@@
is surjective as well, so we can lift the identity map $k\rightarrow k$ to get the desired splitting. To prove the above surjectivity, pick a ring map $h:k\rightarrow R/M^{n}$ where $n\geqslant 1$ is fixed. Let $A:=R/M^{n+1}\times_{R/M^{n}}k$ be the pullback of the natural ring map $\pi:R/M^{n+1}\rightarrow R/M^{n}$ and $h$ in the category of commutative rings. We have then the following commutative diagram of rings:
where $f$ and $f^{\prime}$ are the projection maps. The kernel of $f$ has square-zero. Indeed, it is enough to show that $xy=0$ for all $x=(a+M^{n+1},a^{\prime}+M)$ and $y=(b+M^{n+1},b^{\prime}+M)$ in $\operatorname{Ker}(f)$. We have $f(x)=a^{\prime}+M=0$ and so $a^{\prime}\in M$, but $x\in A$ and so $a+M^{n}=h(a^{\prime}+M)=0$, this shows that $a\in M^{n}$. Similarly, we have $b^{\prime}\in M$ and $b\in M^{n}$. Then $ab\in M^{2n}\subseteq M^{n+1}$ and so $xy=(ab+M^{n+1},a^{\prime}b^{\prime}+M)=0$. The map $f$ is surjective, because the natural ring map $R/M^{n+1}\rightarrow R/M^{n}$ is surjective. In addition, $A$ is a $\mathbb{Q}$-algebra. To see this, it suffices to show that the image of every nonzero integer $d$ under the canonical (injective) ring map $\mathbb{Z}\rightarrow A$ is invertible in $A$. We have $d\cdot 1_{A}=(d\cdot 1_{R}+M^{n+1},d\cdot 1_{R}+M)$. Since $k=R/M$ is a field of characteristic zero, $d\cdot 1_{R}+M$ is invertible in $k$, and so $d\cdot 1_{R}+M^{n}=h(d\cdot 1_{R}+M)$ is invertible in $R/M^{n}$. Then, by the proof of Lemma 2.1, $d\cdot 1_{R}+M^{n+1}\in(R/M^{n+1})^{\ast}$. Thus $d\cdot 1_{A}\in A^{\ast}$. Also, the projection map $f:A\rightarrow k$ is a $\mathbb{Q}$-algebra map, because if $m/d\in\mathbb{Q}$ then $(f\pi_{A})(m/d)=f((m\cdot 1_{A})(d\cdot 1_{A})^{-1})=(m\cdot 1_{R}+M)(d\cdot 1_{R}+M)^{-1}=(m\cdot 1_{k})(d\cdot 1_{k})^{-1}=\pi_{k}(m/d)$ where $\pi_{A}:\mathbb{Q}\rightarrow A$ and $\pi_{k}:\mathbb{Q}\rightarrow k$ are the canonical ring maps. Then by Lemma 3.10, there is a ring map $g:k\rightarrow A$ such that $fg:k\rightarrow k$ is the identity map. Then the composition of the projection map $f^{\prime}:A\rightarrow R/M^{n+1}$ with $g$ gives us the desired ring map $k\rightarrow R/M^{n+1}$.
(ii): First note that as $R$ is complete, the canonical ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ is an isomorphism, and thus the induced map between the groups of units is an isomorphism of groups. We also have $(\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n})^{\ast}=\lim\limits_{\overleftarrow{n\geqslant 1}}(R/M^{n})^{\ast}$. Now to prove the assertion, similar to above, it suffices to show that the natural map:
@@
is surjective for all $n\geqslant 1$ where, as is clear from the context, the $\operatorname{Hom}_{\mathbb{Z}}(-,-)$ is computed in the category of Abelian groups. Suppose we have proven this, then using Remarks 3.2 and 3.15, the natural map:
@@
is surjective as well, so we can lift the identity map $k^{\ast}\rightarrow k^{\ast}$ to get the desired splitting. To prove the above surjectivity, by general homological algebra, it suffices to show that $\operatorname{Ext}^{1}_{\mathbb{Z}}(k^{\ast},A)=0$, where $A=(1+M^{n})/(1+M^{n+1})$. Indeed, using Lemma 3.1, from the following short exact sequence of Abelian groups:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n+1})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$
@@
${\operatorname{Hom}_{\mathbb{Z}}(k^{\ast},(R/M^{n+1})^{\ast})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{\mathbb{Z}}(k^{\ast},(R/M^{n})^{\ast})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}_{\mathbb{Z}}^{1}(k^{\ast},A).}$
But $k^{\ast}$ has trivial $p$-torsion. Also, by Lemma 3.1, the group $A$ is vanished by $p$. Then by Corollary 3.13, we have $\operatorname{Ext}_{\mathbb{Z}}^{1}(k^{\ast},A)=0$. This completes the proof. ∎
@@
This follows from Corollary 3.17. ∎
@@
Under the light of Corollary 3.17, it is natural to ask, is there an example of a local ring $(R,M,k)$ such that the exact sequence of Abelian groups:
@@
is not split? Here we provide such an example. First note that if $M$ is a maximal ideal of a ring $R$ then the following rings $R/M^{n}$ and $R_{M}/M^{n}R_{M}$ are canonically isomorphic for all $n\geqslant 0$. Now let $p$ be a prime number. Consider the local ring $\mathbb{Z}_{P}$ where $P=p\mathbb{Z}$. By the above, $\mathbb{Z}_{P}/P\mathbb{Z}_{P}\simeq\mathbb{Z}/p$. Then the natural surjective map between the groups of units $(\mathbb{Z}_{P})^{\ast}\rightarrow(\mathbb{Z}/p)^{\ast}$ does not admit a splitting for $p\geqslant 5$, because if this map splits then $(\mathbb{Z}/p)^{\ast}$ can be viewed as a subgroup of $(\mathbb{Z}_{P})^{\ast}$, but $\mathbb{Z}_{P}$ does not contain roots of unit of order p1p-1. Thus the above short exact sequence does not split for most incomplete local rings. But for $p=2,3$ the map $(\mathbb{Z}_{P})^{\ast}\rightarrow(\mathbb{Z}/p)^{\ast}$ admits a splitting. We know that for any prime number $p$, the local ring $\mathbb{Z}_{P}$ is not complete (the above observation gives an alternative proof for this fact for $p\geqslant 5$). Thus the above short exact sequence may be split for some incomplete local rings (in other words, the converse of Corollary 3.17 is not true).
The following key result allows us to prove Lemma 1.2, especially in the positive characteristic case.
Theorem 3.20.
@@
Let $k/F$ be an extension of fields with $F$ a perfect field. It is well known that every finitely generated field extension over a perfect field has a separating transcendence basis (for its proof see [7, Theorem 1], [10, Chap. II, §13, Theorem 31] or [8, Theorems 26.2 and 26.3]). This shows that the extension $F\subseteq k$ is separable (in the sense of Definition 3.5), and so by [2, Tag 0320] or by [5, §19, Theorem 19.6.1], this extension is formally smooth. ∎
@@
By Theorem 3.20, the extension $k/F$ is formally smooth. Using this observation and Remark 3.3, we get that the natural map $\operatorname{Hom}_{F}(k,A)\rightarrow\operatorname{Hom}_{F}(k,k)$ given by $g^{\prime}\mapsto fg^{\prime}$ is surjective. In particular, the identity map $k\rightarrow k$ can be lifted to get the desired splitting. ∎
@@
This is a special case of Corollary 3.21. ∎
@@
It is clear that Lemma 3.10 is also a special case of Corollary 3.21. However, the original proof of Lemma 3.10 has the advantage of using less machinery. But the original proof of Lemma 3.10 no longer works to derive Corollary 3.22. In fact, the only key point related with the characteristic used in proving Lemma 3.10 was that every finite extension in characteristic zero is separable. But this does not necessarily hold in the positive characteristic case. For instance, a finite extension of the rational function field $\mathbb{F}_{p}(x)$ is not necessarily separable with $p$ a prime number. For example, the finite extension $\mathbb{F}_{p}(x)\subseteq\mathbb{F}_{p}(x^{1/p})$ is inseparable, because the minimal polynomial of $x^{1/p}$ over $\mathbb{F}_{p}(x)$ has repeated roots: $T^{p}-x=(T-x^{1/p})^{p}$.
@@
Using the above observations, then the following result is the most general version of Theorem 3.16.
@@
is surjective for all $n\geqslant 1$ where the $\operatorname{Hom}(-,-)$ is computed in the category of commutative rings. Take a ring map $h:k\rightarrow R/M^{n}$ where $n\geqslant 1$ is fixed, and let $A:=R/M^{n+1}\times_{R/M^{n}}k$ be the pullback of the natural ring map $R/M^{n+1}\rightarrow R/M^{n}$ and $h$ in the category of commutative rings. Then we show that the projection map $f:A\rightarrow k$ satisfies the conditions of Corollary 3.22. As proved in Theorem 3.16(i), it can be seen that $f$ is surjective and its kernel is of square-zero. We also need to show that $A$ is an $\mathbb{F}_{p}$-algebra. Since $\operatorname{char}(R)=p$, it can be easily seen that $\operatorname{char}(R/M^{d})=p$ for all $d\geqslant 1$. In addition, since the characteristic of a finite direct product of (commutative) rings is indeed the least common multiple (lcm) of the characteristics of the factor rings, and since any extension of rings have the same characteristics, thus $\operatorname{char}(A)=p$, i.e., $A$ is an $\mathbb{F}_{p}$-algebra. Similarly, as proved in Theorem 3.16(i), the projection map $f$ is a morphism of $\mathbb{F}_{p}$-algebras. Then by Corollary 3.22 (or by Corollary 3.21), there is a ring map $g:k\rightarrow A$ such that $fg:k\rightarrow k$ is the identity map. Now the composition of the projection map $A\rightarrow R/M^{n+1}$ with $g$ gives us the desired ring map $k\rightarrow R/M^{n+1}$. This completes the proof. ∎

### v3

@@ 3. The general case
Let $S=\{x_{i}:i\in I\}\subseteq k$ be a transcendence basis of the extension $k/\mathbb{Q}$. Since the set $S$ is algebraically independent over $\mathbb{Q}$, the natural morphism of $\mathbb{Q}$-algebras from the polynomial ring $\mathbb{Q}[X_{i}:i\in I]$ to $k$ given by $X_{i}\mapsto x_{i}$ is injective. Let $R=\mathbb{Q}[x_{i}:i\in I]$ be the image of this ring map. Since $f$ is surjective, for each $i$, there exists some $a_{i}\in A$ such that $f(a_{i})=x_{i}$. Then again by the universal property of polynomial rings, we have a (unique) morphism of $\mathbb{Q}$-algebras $R\rightarrow A$ which maps each $x_{i}$ into $a_{i}$.
To extend this map to $k_{0}=\mathbb{Q}(x_{i}:i\in I)$, the field of fractions of the integral domain $R$, we need to show that the image of every nonzero element $P(x_{i_{1}},\ldots,x_{i_{n}})$ of $R$ under this map is invertible in $A$. Note that $A$ is a local ring with the maximal ideal $M=\operatorname{Ker}(f)$, because $M$ is nilpotent. So it suffices to show that $P(a_{i_{1}},\ldots,a_{i_{n}})\notin M$. Since $f$ is a $\mathbb{Q}$-algebra map, we have $f(P(a_{i_{1}},\ldots,a_{i_{n}}))=P(x_{i_{1}},\ldots,x_{i_{n}})\neq 0$. This shows that $P(a_{i_{1}},\ldots,a_{i_{n}})$ is invertible in $A$. So we can extend the above ring map to obtain a ring map $h:k_{0}\rightarrow A$ such that $fh:k_{0}\rightarrow k$ is the inclusion map. This, in particular, shows that $A$ is a $k_{0}$-algebra and $h$ is a $k_{0}$-algebra morphism.
To extend this ring map $h$ to $k$, since $k$ is an algebraic extension over $k_{0}$, then by Corollary 3.9, the extension $k/k_{0}$ is formally étale. This observation allows us to show that the natural map $\operatorname{Hom}_{k_{0}}(k,A)\rightarrow\operatorname{Hom}_{k_{0}}(k,k)$ given by $g^{\prime}\mapsto fg^{\prime}$ is bijective where the $\operatorname{Hom}_{k_{0}}(-,-)$ is computed in the category of $k_{0}$-algebras. Indeed, if $h^{\prime}:k\rightarrow k$ is a morphism of $k_{0}$-algebras then we have the following commutative diagram of rings:
@@
The key point in the preceding proof is formally smoothness of $k$ over $\mathbb{Q}$, which Lemma 3.19 (and Corollary 3.21) will adapt to any perfect field in place of $\mathbb{Q}$.
@@
Consider the following short exact sequence of $R$-modules:
@@
where the second map is multiplication by $r$ on the $\operatorname{Ext}_{R}^{d}(M,N)$. But by [2, Tag 00LV], this map vanishes since $rN=0$, then in either case of the hypothesis, making $\operatorname{Ext}_{R}^{d}(M,N)$ vanish by exactness. ∎
@@
The above corollary is quite sufficient for our purpose. However, we should add that this result holds more generally: Let $R$ be a regular Noetherian ring of Krull dimension $n$ and $M,N$ be modules over $R$. If there is an element $r\in R$ such that $r$ is regular on $M$ and $rN=0$, then $\operatorname{Ext}_{R}^{n}(M,N)=0$. In fact, this is proved by the same argument as the above by applying a theorem of Serre [9] which asserts that if $R$ is regular Noetherian of Krull dimension $n$, then for any two $R$-modules $M,N$ one has $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>n$. The above corollary also holds over a Dedekind domain as well.
@@
$\mathbf{(ii)}$ If $\operatorname{char}(k)=p>0$, then the natural surjective group map $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
@@
given by $g\mapsto\pi g$ is surjective for all $n\geqslant 1$, where $\pi:R/M^{n+1}\rightarrow R/M^{n}$ is the natural ring map and the $\operatorname{Hom}(-,-)$ is computed in the category of commutative rings. Suppose we have proven this, then using Remarks 3.2 and 3.14, the natural map:
@@
given by $g\mapsto\pi^{\prime}g$ is surjective as well where $\pi^{\prime}:R\rightarrow k$ is the natural ring map, so we can lift the identity map $k\rightarrow k$ to get the desired splitting.
To prove the above surjectivity, pick a ring map $h:k\rightarrow R/M^{n}$ where $n\geqslant 1$ is fixed. Let $A:=R/M^{n+1}\times_{R/M^{n}}k$ be the pullback of the natural ring map $\pi:R/M^{n+1}\rightarrow R/M^{n}$ and $h$ in the category of commutative rings. We have then the following commutative diagram of rings:
where $f$ and $f^{\prime}$ are the projection maps. The kernel of $f$ has square-zero. Indeed, it is enough to show that $xy=0$ for all $x=(a+M^{n+1},a^{\prime}+M)$ and $y=(b+M^{n+1},b^{\prime}+M)$ in $\operatorname{Ker}(f)$. We have $f(x)=a^{\prime}+M=0$ and so $a^{\prime}\in M$, but $x\in A$ and so $a+M^{n}=h(a^{\prime}+M)=0$, this shows that $a\in M^{n}$. Similarly, we have $b^{\prime}\in M$ and $b\in M^{n}$. Then $ab\in M^{2n}\subseteq M^{n+1}$ and so $xy=(ab+M^{n+1},a^{\prime}b^{\prime}+M)=0$.
The map $f$ is surjective, because the natural ring map $R/M^{n+1}\rightarrow R/M^{n}$ is surjective.
In addition, $A$ is a $\mathbb{Q}$-algebra. To see this, it suffices to show that the image of every nonzero integer $d$ under the natural (injective) ring map $\mathbb{Z}\rightarrow A$ is invertible in $A$. We have $d\cdot 1_{A}=(d\cdot 1_{R}+M^{n+1},d\cdot 1_{R}+M)$. Since $k=R/M$ is a field of characteristic zero, $d\cdot 1_{R}+M$ is invertible in $k$, and so $d\cdot 1_{R}+M^{n}=h(d\cdot 1_{R}+M)$ is invertible in $R/M^{n}$. Then, by the proof of Lemma 2.1, $d\cdot 1_{R}+M^{n+1}\in(R/M^{n+1})^{\ast}$. Thus $d\cdot 1_{A}\in A^{\ast}$.
Also, the projection map $f:A\rightarrow k$ is a $\mathbb{Q}$-algebra map, because if $m/d\in\mathbb{Q}$ then $(f\pi_{A})(m/d)=f((m\cdot 1_{A})(d\cdot 1_{A})^{-1})=(m\cdot 1_{R}+M)(d\cdot 1_{R}+M)^{-1}=(m\cdot 1_{k})(d\cdot 1_{k})^{-1}=\pi_{k}(m/d)$ where $\pi_{A}:\mathbb{Q}\rightarrow A$ and $\pi_{k}:\mathbb{Q}\rightarrow k$ are the natural ring maps.
Then by Lemma 3.10, there is a ring map $g:k\rightarrow A$ such that $fg:k\rightarrow k$ is the identity map. Then the composition of the projection map $f^{\prime}:A\rightarrow R/M^{n+1}$ with $g$ gives us the desired ring map $k\rightarrow R/M^{n+1}$.
(ii): First note that as $R$ is complete, the natural ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ is an isomorphism, and thus the induced map between the groups of units is an isomorphism of groups. We also have $(\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n})^{\ast}=\lim\limits_{\overleftarrow{n\geqslant 1}}(R/M^{n})^{\ast}$. Now to prove the assertion, similar to above, it suffices to show that the natural map:
@@
given by $g\mapsto\pi^{\ast}g$ is surjective for all $n\geqslant 1$ where $\pi:R/M^{n+1}\rightarrow R/M^{n}$ is the natural ring map and the $\operatorname{Hom}_{\mathbb{Z}}(-,-)$ is computed in the category of Abelian groups. Suppose we have proven this, then using Remarks 3.2 and 3.14, the natural map:
@@
is surjective as well, so we can lift the identity map $k^{\ast}\rightarrow k^{\ast}$ to get the desired splitting. To prove the above surjectivity, by general homological algebra, it suffices to show that $\operatorname{Ext}^{1}_{\mathbb{Z}}(k^{\ast},G)=0$, where $G=(1+M^{n})/(1+M^{n+1})$. Indeed, using Lemma 3.1, from the following short exact sequence of Abelian groups:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n+1})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$
@@
${\operatorname{Hom}_{\mathbb{Z}}(k^{\ast},(R/M^{n+1})^{\ast})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{\mathbb{Z}}(k^{\ast},(R/M^{n})^{\ast})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}_{\mathbb{Z}}^{1}(k^{\ast},G).}$
But $k^{\ast}$ has trivial $p$-torsion. Also, by Lemma 3.1, the Abelian group $G$ is vanished by $p$. Then by Corollary 3.12, we have $\operatorname{Ext}_{\mathbb{Z}}^{1}(k^{\ast},G)=0$. This completes the proof. ∎
@@
This follows from Corollary 3.16 and Remark 2.5. ∎
@@
We give an example of an (incomplete) local ring $(R,M,k)$ such that the following exact sequence of Abelian groups:
@@
is not split. First note that if $M$ is a maximal ideal of a ring $R$ then the rings $R/M^{n}$ and $R_{M}/M^{n}R_{M}$ are naturally isomorphic for all $n\geqslant 0$.
Now let $p$ be a prime number. Consider the local ring $\mathbb{Z}_{P}$ where $P=p\mathbb{Z}$. By the above, $\mathbb{Z}_{P}/P\mathbb{Z}_{P}\simeq\mathbb{F}_{p}=\mathbb{Z}/p$. Then the natural surjective map between the groups of units $\mathbb{Z}_{P}^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ does not admit a splitting for $p\geqslant 5$, because if this map splits then $\mathbb{F}_{p}^{\ast}$ can be viewed as a subgroup of $\mathbb{Z}_{P}^{\ast}$, but $\mathbb{Z}_{P}$ does not contain roots of unit of order p1p-1.
But for $p=2,3$ the map $\mathbb{Z}_{P}^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ admits a splitting. We know that for any prime number $p$, the local ring $\mathbb{Z}_{P}$ is not complete. This shows that the above short exact sequence may be split for some incomplete local rings (in other words, the converse of Corollary 3.16 is not true).
The following result allows us to prove Lemma 1.2, especially in the positive characteristic case.
Lemma 3.19.
@@
Let $k/F$ be an extension of fields with $F$ a perfect field. It is well known that every finitely generated field extension over a perfect field has a separating transcendence basis (for its proof see [7, Theorem 1] or [10, Chap. II, §13, Theorem 31]). This shows that the extension $F\subseteq k$ is separable (in the sense of Definition 3.5), and so by [2, Tag 0320] or by [5, §19, Theorem 19.6.1], this extension is formally smooth. As a second proof, combine Theorems 26.3 and 26.9 from Matsumura’s book [8]. ∎
Corollary 3.20.
In the characteristic zero, every extension of fields is formally smooth.
Proof.
Every field of characteristic zero is perfect. Then apply Lemma 3.19. ∎
@@
By Lemma 3.19, the extension $k/F$ is formally smooth. Using this observation and Remark 3.3, we get that the natural map $\operatorname{Hom}_{F}(k,A)\rightarrow\operatorname{Hom}_{F}(k,k)$ given by $g^{\prime}\mapsto fg^{\prime}$ is surjective where the $\operatorname{Hom}_{F}(-,-)$ is computed in the category of $F$-algebras. In particular, the identity map $k\rightarrow k$ can be lifted to get the desired splitting. ∎
@@
This is a special case of Corollary 3.21 (because every finite field is a perfect field). ∎
@@
It is clear that Lemma 3.10 is also a special case of Corollary 3.21. However, the original proof of Lemma 3.10 has the important advantage of using less machinery. Also note that the original proof of Lemma 3.10 no longer works to derive Corollary 3.22. In fact, the only key point related with the characteristic used in proving Lemma 3.10 was that, in the characteristic zero, every algebraic extension of fields is separable. But this does not necessarily hold in the positive characteristic case. For instance, a finite extension of the rational function field $\mathbb{F}_{p}(x)$ is not necessarily separable with $p$ a prime number. For example, the finite extension $\mathbb{F}_{p}(x)\subseteq\mathbb{F}_{p}(x^{1/p})$ is inseparable, because the minimal polynomial of $x^{1/p}$ over $\mathbb{F}_{p}(x)$ has repeated roots: $T^{p}-x=(T-x^{1/p})^{p}$.
@@
The following result is the most general version of Theorem 3.15.
@@
is surjective for all $n\geqslant 1$ where the $\operatorname{Hom}(-,-)$ is computed in the category of commutative rings.
Take a ring map $h:k\rightarrow R/M^{n}$ where $n\geqslant 1$ is fixed, and let $A:=R/M^{n+1}\times_{R/M^{n}}k$ be the pullback of the natural ring map $R/M^{n+1}\rightarrow R/M^{n}$ and $h$ in the category of commutative rings. Then we show that the projection map $f:A\rightarrow k$ satisfies the conditions of Corollary 3.22. As proved in Theorem 3.15(i), it can be seen that $f$ is surjective and its kernel is of square-zero.
We also need to show that $A$ is an $\mathbb{F}_{p}$-algebra. Since $\operatorname{char}(R)=p$, it can be easily seen that $\operatorname{char}(R/M^{d})=p$ for all $d\geqslant 1$. In addition, since the characteristic of a finite direct product of (commutative) rings is indeed the least common multiple (lcm) of the characteristics of the factor rings, and since any extension of rings have the same characteristics, thus $\operatorname{char}(A)=p$, i.e., $A$ is an $\mathbb{F}_{p}$-algebra. Similarly, as proved in Theorem 3.15(i), the projection map $f$ is a morphism of $\mathbb{F}_{p}$-algebras.
Then by Corollary 3.22 (or by Corollary 3.21), there is a ring map $g:k\rightarrow A$ such that $fg:k\rightarrow k$ is the identity map. Now the composition of the projection map $A\rightarrow R/M^{n+1}$ with $g$ gives us the desired ring map $k\rightarrow R/M^{n+1}$. This completes the proof. ∎